OpenStudy (anonymous):

NaHCO3(s)+HCl(aq)-->NaCl(aq)+ H2O can someone show me how to solve this

OpenStudy (anonymous):

Where did your CO2 go?Anyway this is maths section

OpenStudy (anonymous):

but can you help me on chem then

OpenStudy (anonymous):

@Nishant_Garg

OpenStudy (anonymous):

What do you need help with?I didn't exactly understand your question

OpenStudy (anonymous):

do you know how to do stoicheometry

OpenStudy (anonymous):

@Nishant_Garg

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@Nishant_Garg

OpenStudy (anonymous):

So what exactly is your question?

OpenStudy (anonymous):

NaHCO3(s)+HCl(aq)-->NaCl(aq)+ H2O how to get grams to moles to moles to grams

OpenStudy (anonymous):

Calculate the molecular mass of a molecule, 1 mole of that molecule is then equal to the molar mass express in grams The molecular mass of NaHCO3 is \[23+1+12+(16 \times 3)=36+48=84\] Thus 84g of NaHCO3 is equal to 1 mol of NaHCO3

OpenStudy (anonymous):

similarly 1 mol of H atoms is equal to 1g of H atom, 1 mol of Sodium atoms is equal to 23g of sodium

OpenStudy (anonymous):

note that 1 mol of hydrogen gas is 2grams because when we talk about hydrogen gas that's \[H_{2}\], that's 2 atoms so its 2grams

OpenStudy (anonymous):

thank you sooo much

OpenStudy (anonymous):

Similarly oxygen GAS/molecule\[O_{2}\] 1 mole is 32g but 1 mol of oxygen ATOMS is 16g